A charge of 9.00 mC is placed at each corner of a square 0.270 m on a side.

Determine the magnitude of the force on each charge.
Answer #1

given charge q = 9 m C = 0.009 C

r = 0.27 m

F1 = F4 = k q1 q2 / r2   

= 9 X 109 X 0.0092 / 0.272

= 107 N

the distance is a = ( 0.272 + 0.272 )1/2

a = 0.381 m

F3 = k q3 q2 / a2

= 9 X 109 X 0.0092 / 0.3812

= 5.022 X 106 N

the angle is \theta = 450

X - component is

Fx = F1 + F3 sin\theta

= 107 + 5.022 X 106 X sin45

= 1.355 X 107 N

Fy = F4 + F3 sin\theta

= 107 + 5.022 X 106 X sin45

= 1.355 X 107 N

F = ( Fx2 + Fy2 )1/2

F = ( (1.355 X 107 )2 + (1.355 X 107 )2 )1/2

F = 19127271.11 N

the magnitude of the force on each charge is F = 19127271.11 N